# Operator World Models for Reinforcement Learning

### v1

@@ A.3 The operatorial formulation of RL
$q({P},{T}_{1},r_{1})-q({P},{T}_{2},r_{2})$ $=({Id}-\gamma{T}_{1}{P})^{-1}r_{1}-({Id}-\gamma{T}_{2}{P})^{-1}r_{2}$
$=({Id}-\gamma{T}_{1}{P})^{-1}(r_{1}-r_{2})+\left[({Id}-\gamma{T}_{1}{P})^{-1}-({Id}-\gamma{T}_{2}{P})^{-1}\right]r_{2}$
where we added and removed a term $({Id}-\gamma{T}_{1}{P})^{-1}r_{2}$. The result follows by plugging in the Simulation Lemma A.7 for the second term of the right hand side. ∎
The corollary above will be useful in Appendix C to control the approximation error of the estimates $\hat{q}_{t}$ appearing in the convergence rates for inexact PMD in Thm. 7.
@@ A.4 Action-value Estimator for $({\mathcal{G}},{\mathcal{F}})$-compatible Policies
$\hat{q}_{\pi}=({Id}-\gamma{T}_{n}{P}_{\pi})^{-1}r_{n}=({Id}-\gamma S_{n}^{*}BZ_{n}{P}_{\pi})^{-1}S_{n}^{*}b.$ (A.14)
@@
$(M_{\pi n})_{ij}=\left\langle{e_{i}},{M_{\pi n}e_{j}}\right\rangle=\left\langle{e_{i}},{Z_{n}{P}_{\pi}S_{n}^{*}}\right\rangle=\left\langle{Z_{n}^{*}e_{i}},{{P}_{\pi}e_{j}}\right\rangle.$ (A.15)
Now, we recall that the two operators $S_{n}:{\mathcal{G}}\to\mathbb{R}^{n}$ and $Z_{n}^{*}:{\mathcal{F}}\to\mathbb{R}^{n}$ are the evaluation operators for respectively the points $(x_{i},a_{i})_{i=1}^{n}$ and $(x_{i}^{\prime})_{i=1}^{n}$. Namely, for any vector $v\in\mathbb{R}^{n}$
@@
$(M_{\pi n})_{ij)}=\left\langle{Z_{n}^{*}e_{i}},{{P}_{\pi}S_{n}^{*}e_{j}}\right\rangle=\left\langle{\varphi(x_{i}^{\prime})},{{P}_{\pi}\psi(x_{j},a_{j})}\right\rangle$ (A.17)
Since ${P}_{\pi}$ is $({\mathcal{G}},{\mathcal{F}})$-compatible by hypothesis, we can leverage the same reasoning used in Prop. 2 to show that
$({P}_{\pi}|_{\mathcal{G}})^{*}\varphi(x^{\prime})=\int_{\mathcal{A}}\psi(x^{\prime},a)\leavevmode\nobreak\ \pi(da|x^{\prime})$ (A.18)
for any $x^{\prime}\in\mathcal{X}$. By plugging this equation in the previous characterization for $(M_{\pi n})_{ij}$ we have
$\left\langle{\varphi(x_{i}^{\prime})},{{P}_{\pi}\psi(x_{j},a_{j})}\right\rangle$ $=\left\langle{{P}_{\pi}^{*}\varphi(x_{i}^{\prime})},{\psi(x_{j},a_{j})}\right\rangle$ (A.19)
@@ A.5 Separable Spaces
The proposition follows from observing that for any $v\in\mathbb{R}^{|\mathcal{A}|}$ and $h\in{\mathcal{H}}$, applying ${P}_{\pi}$ according to 6 to the function $g(x,a)=\left\langle{h},{\phi(x)}\right\rangle\left\langle{v},{De_{a}}\right\rangle$ yields
$({P}_{\pi}g)(x)=\sum_{a\in\mathcal{A}}g(x,a)\pi(a|x)$ $=\left\langle{h},{\phi(x)}\right\rangle\sum_{a\in\mathcal{A}}\left\langle{v},{De_{a}}\right\rangle\pi(a|x)$ (A.22)
@@
Hence $({P}_{\pi}g)(x)=\left\langle{f},{\varphi(x)}\right\rangle$ with $f=h\otimes h^{\prime}\in{\mathcal{H}}\otimes{\mathcal{H}}={\mathcal{F}}$ and $h^{\prime}=\sum_{a\in\mathcal{A}}\left\langle{v},{De_{a}}\right\rangle p_{a}\in{\mathcal{H}}$. Therefore, the restriction of ${P}_{\pi}$ to ${\mathcal{G}}$ takes value in ${\mathcal{F}}$ as desired. ∎
@@ Appendix B Policy Mirror Descent
Let $D:\Delta(\mathcal{A})\times\text{rint}\Delta(\mathcal{A})\to\mathbb{R}$ a Bregman divergence [31, Definition 9.2] over the probability simplex, where $\text{rint}\Delta(\mathcal{A})$ denotes the relative interior of $\Delta(\mathcal{A})$. In the following, for any $t\in\mathbb{N}$ we will denote by $\pi_{t}$ the policy produced at iteration $t$ by a PMD algorithm according to the update 2 (with either the exact action-value function or an estimator, as discussed in Sec. 3) with divergence $D$ and step-size $\eta>0$. We denote ${P}_{t}={P}_{\pi_{t}}$ the associated operator. We recall here the PMD update from 2, highlighting the dependency on the policy operator ${P}_{t}$ via the action-value function $q({P}_{t})$.
$\pi_{t+1}(\cdot|x)\in\operatorname*{argmin}_{p\in\Delta(\mathcal{A})}\left\{-\eta\sum_{a\in\mathcal{A}}q({P}_{t})(x,a)p_{a}+D(p;\pi_{t}(\cdot|x))\right\}\quad\text{for all }x\in\mathcal{X}.$ (B.1)
While this point-wise characterization is sufficient to define the updated policy $\pi_{t+1}:\mathcal{X}\to\Delta(\mathcal{A})$ from the previous $\pi_{t}$ and its action-value function $q({P}_{t})$, we need to guarantee that $\pi_{t+1}$ is measurable. If that were not the case, we would not be able to guarantee the existence of a ${P}_{t+1}$ associated with it, possibly affecting the well-definiteness of iteratively applying the mirror descent update B.1. The following result addresses this issue.
@@
The proof follows from the Measurable Maximum Theorem [25, Theorem 18.19]. Let us denote $f_{t}:\mathcal{X}\times\Delta(\mathcal{A})\to\mathbb{R}$ the function
$f_{t}(x,p):=-\eta\sum_{a\in\mathcal{A}}q({P}_{t})(x,a)p_{a}+D(p;\pi_{t}(x)).$ (B.2)
Let also $\kappa:\mathcal{X}\twoheadrightarrow\Delta(\mathcal{A})$ be the constant correspondance $x\mapsto\Delta(\mathcal{A})$ for all $x\in\mathcal{X}$. $\kappa$ clearly has nonempty compact values, and it is also weakly measurable since for any open set $G\subset\Delta(\mathcal{A})$, its lower inverse $\kappa^{\ell}(G):=\{x\in\mathcal{X}:\kappa(x)\cap G\neq\varnothing\}=\mathcal{X}$ belongs to the Borel sigma-algebra of $\mathcal{X}$. Finally, since $q({P}_{t})\in B_{b}(\Omega)$, and by assumption $D$ is continuous in its first argument, then we have that $f_{t}$ is a Carathéodory function. Then, by [25, Theorem 18.19] we have that the correspondance of minimizers $\mu:\mathcal{X}\twoheadrightarrow\Delta(\mathcal{A})$ defined as
@@
The previous Lemma is the key technical step enabling us to extend the convergence rates of Mirror Descent proved in [12] to non-tabular settings. We now state and prove few Lemmas instrumental to prove Thm. 7.
@@
Let $\pi_{t+1}:\mathcal{X}\to\Delta(\mathcal{A})$ a measurable minimizer of B.2 and ${P}_{t+1}$ its associated operator. For every measurable policy $\pi:\mathcal{X}\to\Delta(\mathcal{A})$ (alongside its associated operator ${P}$) it holds
$\eta\left[({P}_{t+1}-{P})q({P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))$ (B.3)
@@
The function $f_{t}(x,p)$ in B.2 is convex and differentiable in $p$ as it is a sum of a linear function and a (strictly convex) Bregman divergence. By the first-order optimality condition [31, Corollay 3.68], a minimizer $p_{*}$ of $f_{t}(x,\cdot)$ satisfies, for all $p\in\Delta(\mathcal{A})$
@@
$\begin{split}&\eta\left[({P}_{t+1}-{P})q({P}_{t})\right](x)-\left\langle{\nabla\psi(\pi_{t}(x))-\nabla\psi(\pi_{t+1}(x))},{\pi(x)-\pi_{t+1}(x)}\right\rangle\geq 0\quad\implies\\ &\eta\left[({P}_{t+1}-{P})q({P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))+D(\pi_{t+1}(x);\pi_{t}(x))\quad\implies\\ &\eta\left[({P}_{t+1}-{P})q({P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))\end{split}$
Where in the first line we used the defintion of Bregman divergence [31, Definition 9.2] $D(p;q):=\psi(p)-\psi(q)-\left\langle{\nabla\psi(q)},{p-q}\right\rangle$ for a suitable Legendre function $\psi:\Delta(\mathcal{A})\to\mathbb{R}$, the first implication follows from the three-points property of Bregman divergences [31, Lemma 9.11], and the last implication from the positivity of $D(\pi_{t+1}(x);\pi_{t}(x))$. ∎
@@
This Corollary is essentially a restatement of [12, Lemma 7]. Let $({P}_{t})_{t\in\mathbb{N}}$ be the sequence of policy operators associated to the measurable minimizers of B.1 for all $t\in\mathbb{N}$. For all $x\in\mathcal{X}$ it holds
$\left[({P}_{t+1}-{P}_{t})q({P}_{t})\right](x)\geq 0$ (B.5)
@@
$J({P}_{t+1})-J({P}_{t})\geq 0$ (B.6)
i.e. the objective function is always increased by a mirror descent iteration. Further, if $\tilde{q}({P}_{t})\in B_{b}(\Omega)$ is such that $\left\|{q({P}_{t})-\tilde{q}({P}_{t})}\right\|_{\infty}\leq\varepsilon_{t}$, then B.5 holds inexactly on $\tilde{q}({P}_{t})$ as
$\left[({P}_{t+1}-{P}_{t})\tilde{q}({P}_{t})\right](x)\geq-2\varepsilon_{t}.$ (B.7)
@@
$\eta\left[({P}_{t+1}-{P}_{t})q({P}_{t})\right](x)\geq D(\pi_{t}(x);\pi_{t+1}(x))\geq 0,$
giving B.5. Integrating B.5 over $\left({Id}-\gamma{P}_{t+1}{T}\right)^{-*}\nu$ and using the Performance Difference Lemma A.4 one gets B.6. Finally, we get B.7 from
$\begin{split}\left[({P}_{t+1}-{P}_{t})\tilde{q}({P}_{t})\right](x)&=\left[({P}_{t+1}-{P}_{t})q({P}_{t})\right](x)+\left[({P}_{t+1}-{P}_{t})(\tilde{q}({P}_{t})-q({P}_{t}))\right](x)\\ &\geq\left[({P}_{t+1}-{P}_{t})(\tilde{q}({P}_{t})-q({P}_{t}))\right](x)\\ &\geq-\bigl{\|}{({P}_{t+1}-{P}_{t})(\tilde{q}({P}_{t})-q({P}_{t}))}\bigr{\|}_{\infty}\\ &\geq-\bigl{\|}{{P}_{t+1}-{P}_{t}}\bigr{\|}\bigl{\|}{\tilde{q}({P}_{t})-q({P}_{t})}\bigr{\|}_{\infty}\\ &\geq-2\varepsilon_{t}.\end{split}$ (B.8)
Where the first inequality follows from B.5, and the latter from the fact that policy operators are Markov operators and have norm 1, and $\bigl{\|}{{P}_{t+1}-{P}_{t}}\bigr{\|}\leq\bigl{\|}{{P}_{t+1}}\bigr{\|}+\bigl{\|}{{P}_{t}}\bigr{\|}=2$. ∎

### v2

@@ A.3 The operatorial formulation of RL
$q(\mathsf{P},\mathsf{T}_{1},r_{1})-q(\mathsf{P},\mathsf{T}_{2},r_{2})$ $=(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}r_{1}-(\mathsf{Id}-\gamma\mathsf{T}_{2}\mathsf{P})^{-1}r_{2}$
$=(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}(r_{1}-r_{2})+\left[(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}-(\mathsf{Id}-\gamma\mathsf{T}_{2}\mathsf{P})^{-1}\right]r_{2}$
where we added and removed a term $(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}r_{2}$. The result follows by plugging in the Simulation Lemma A.6 for the second term of the right hand side. ∎
The corollary above will be useful in Appendix C to control the approximation error of the estimates $\hat{q}_{\pi_{t}}$ appearing in the convergence rates for inexact PMD in Thm. 7.
@@ A.4 Action-value Estimator for $({\mathcal{G}},{\mathcal{F}})$-compatible Policies
$\hat{q}_{\pi}=(\mathsf{Id}-\gamma\mathsf{T}_{n}\mathsf{P}_{\pi})^{-1}r_{n}=(\mathsf{Id}-\gamma S_{n}^{*}BZ_{n}\mathsf{P}_{\pi})^{-1}S_{n}^{*}b.$ (A.14)
@@
$(M_{\pi})_{ij}=\left\langle{e_{i}},{M_{\pi}e_{j}}\right\rangle=\left\langle{e_{i}},{Z_{n}\mathsf{P}_{\pi}S_{n}^{*}e_{j}}\right\rangle=\left\langle{Z_{n}^{*}e_{i}},{\mathsf{P}_{\pi}S_{n}^{*}e_{j}}\right\rangle.$ (A.15)
Now, we recall that the two operators $S_{n}:{\mathcal{G}}\to{\mathbb{R}}^{n}$ and $Z_{n}:{\mathcal{F}}\to{\mathbb{R}}^{n}$ are the evaluation operators for respectively the points $(x_{i},a_{i})_{i=1}^{n}$ and $(x_{i}^{\prime})_{i=1}^{n}$. Namely, for any vector $v\in{\mathbb{R}}^{n}$
@@
$(M_{\pi})_{ij}=\left\langle{Z_{n}^{*}e_{i}},{\mathsf{P}_{\pi}S_{n}^{*}e_{j}}\right\rangle=\left\langle{\varphi(x_{i}^{\prime})},{\mathsf{P}_{\pi}\psi(x_{j},a_{j})}\right\rangle$ (A.17)
Since $\mathsf{P}_{\pi}$ is $({\mathcal{G}},{\mathcal{F}})$-compatible by hypothesis, we can leverage the same reasoning used in Prop. 2 to show that
$(\mathsf{P}_{\pi}|_{\mathcal{G}})^{*}\varphi(x^{\prime})=\int_{\mathcal{A}}\psi(x^{\prime},a)\leavevmode\nobreak\ \pi(da|x^{\prime})$ (A.18)
for any $x^{\prime}\in\mathcal{X}$. By plugging this equation in the previous characterization for $(M_{\pi})_{ij}$ we have
$\left\langle{\varphi(x_{i}^{\prime})},{\mathsf{P}_{\pi}\psi(x_{j},a_{j})}\right\rangle$ $=\left\langle{\mathsf{P}_{\pi}^{*}\varphi(x_{i}^{\prime})},{\psi(x_{j},a_{j})}\right\rangle$ (A.19)
@@ A.5 Separable Spaces
The proposition follows from observing that for any $v\in{\mathbb{R}}^{|\mathcal{A}|}$ and $h\in{\mathcal{H}}$, applying $\mathsf{P}_{\pi}$ according to 6 to the function $g(x,a)=\left\langle{h},{\phi(x)}\right\rangle\left\langle{v},{De_{a}}\right\rangle$ yields
$(\mathsf{P}_{\pi}g)(x)=\sum_{a\in\mathcal{A}}g(x,a)\pi(a|x)$ $=\left\langle{h},{\phi(x)}\right\rangle\sum_{a\in\mathcal{A}}\left\langle{v},{De_{a}}\right\rangle\pi(a|x)$ (A.22)
@@
Hence $(\mathsf{P}_{\pi}g)(x)=\left\langle{f},{\varphi(x)}\right\rangle$ with $f=h\otimes h^{\prime}\in{\mathcal{H}}\otimes{\mathcal{H}}={\mathcal{F}}$ and $h^{\prime}=\sum_{a\in\mathcal{A}}\left\langle{v},{De_{a}}\right\rangle p_{a}\in{\mathcal{H}}$. Therefore, the restriction of $\mathsf{P}_{\pi}$ to ${\mathcal{G}}$ takes value in ${\mathcal{F}}$ as desired. ∎
@@ Appendix B Policy Mirror Descent
Let $D:\Delta(\mathcal{A})\times\text{rint}\Delta(\mathcal{A})\to{\mathbb{R}}$ a Bregman divergence [35, Definition 9.2] over the probability simplex, where $\text{rint}\Delta(\mathcal{A})$ denotes the relative interior of $\Delta(\mathcal{A})$. In the following, for any $t\in{\mathbb{N}}$ we will denote by $\pi_{t}$ the policy produced at iteration $t$ by a PMD algorithm according to the update 2 (with either the exact action-value function or an estimator, as discussed in Sec. 3) with divergence $D$ and step-size $\eta>0$. We denote $\mathsf{P}_{t}=\mathsf{P}_{\pi_{t}}$ the associated operator. We recall here the PMD update from 2, highlighting the dependency on the policy operator $\mathsf{P}_{t}$ via the action-value function $q(\mathsf{P}_{t})$.
$\pi_{t+1}(\cdot|x)\in\operatornamewithlimits{argmin}_{p\in\Delta(\mathcal{A})}\left\{-\eta\sum_{a\in\mathcal{A}}q(\mathsf{P}_{t})(x,a)p_{a}+D(p;\pi_{t}(\cdot|x))\right\}\quad\text{for all }x\in\mathcal{X}.$ (B.1)
While this point-wise characterization is sufficient to define the updated policy $\pi_{t+1}:\mathcal{X}\to\Delta(\mathcal{A})$ from the previous $\pi_{t}$ and its action-value function $q(\mathsf{P}_{t})$, we need to guarantee that $\pi_{t+1}$ is measurable. If that were not the case, we would not be able to guarantee the existence of a $\mathsf{P}_{t+1}$ associated with it, possibly affecting the well-definiteness of iteratively applying the mirror descent update B.1. The following result addresses this issue.
@@
The proof follows from the Measurable Maximum Theorem [29, Theorem 18.19]. Let us denote $f_{t}:\mathcal{X}\times\Delta(\mathcal{A})\to{\mathbb{R}}$ the function
$f_{t}(x,p):=-\eta\sum_{a\in\mathcal{A}}q(\mathsf{P}_{t})(x,a)p_{a}+D(p;\pi_{t}(x)).$ (B.2)
Let also $\kappa:\mathcal{X}\twoheadrightarrow\Delta(\mathcal{A})$ be the constant correspondance $x\mapsto\Delta(\mathcal{A})$ for all $x\in\mathcal{X}$. $\kappa$ clearly has nonempty compact values, and it is also weakly measurable since for any open set $G\subset\Delta(\mathcal{A})$, its lower inverse $\kappa^{\ell}(G):=\{x\in\mathcal{X}:\kappa(x)\cap G\neq\varnothing\}=\mathcal{X}$ belongs to the Borel sigma-algebra of $\mathcal{X}$. Finally, since $q(\mathsf{P}_{t})\in B_{b}(\Omega)$, and by assumption $D$ is continuous in its first argument, then we have that $f_{t}$ is a Carathéodory function. Then, by [29, Theorem 18.19] we have that the correspondence of minimizers $\mu:\mathcal{X}\twoheadrightarrow\Delta(\mathcal{A})$ defined as
@@
The previous Lemma is the key technical step enabling us to extend the convergence rates of Mirror Descent proved in [12] to non-tabular settings. We now state and prove fa ew Lemmas instrumental to prove Thm. 7.
@@
Let $\pi_{t+1}:\mathcal{X}\to\Delta(\mathcal{A})$ a measurable minimizer of B.2 and $\mathsf{P}_{t+1}$ its associated operator. For every measurable policy $\pi:\mathcal{X}\to\Delta(\mathcal{A})$ (alongside its associated operator $\mathsf{P}$) it holds
$\eta\left[(\mathsf{P}_{t+1}-\mathsf{P})q(\mathsf{P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))$ (B.3)
@@
The function $f_{t}(x,p)$ in B.2 is convex and differentiable in $p$ as it is a sum of a linear function and a (strictly convex) Bregman divergence. By the first-order optimality condition [35, Corollay 3.68], a minimizer $p_{*}$ of $f_{t}(x,\cdot)$ satisfies, for all $p\in\Delta(\mathcal{A})$
@@
$\begin{split}&\eta\left[(\mathsf{P}_{t+1}-\mathsf{P})q(\mathsf{P}_{t})\right](x)-\left\langle{\nabla\psi(\pi_{t}(x))-\nabla\psi(\pi_{t+1}(x))},{\pi(x)-\pi_{t+1}(x)}\right\rangle\geq 0\quad\implies\\ &\eta\left[(\mathsf{P}_{t+1}-\mathsf{P})q(\mathsf{P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))+D(\pi_{t+1}(x);\pi_{t}(x))\quad\implies\\ &\eta\left[(\mathsf{P}_{t+1}-\mathsf{P})q(\mathsf{P}_{t})\right](x)\geq D(\pi(x);\pi_{t+1}(x))-D(\pi(x);\pi_{t}(x))\end{split}$
Where in the first line we used the definition of Bregman divergence [35, Definition 9.2] $D(p;q):=\psi(p)-\psi(q)-\left\langle{\nabla\psi(q)},{p-q}\right\rangle$ for a suitable Legendre function $\psi:\Delta(\mathcal{A})\to{\mathbb{R}}$, the first implication follows from the three-points property of Bregman divergences [35, Lemma 9.11], and the last implication from the positivity of $D(\pi_{t+1}(x);\pi_{t}(x))$. ∎
@@
This Corollary is essentially a restatement of [12, Lemma 7]. Let $(\mathsf{P}_{t})_{t\in{\mathbb{N}}}$ be the sequence of policy operators associated to the measurable minimizers of B.1 for all $t\in{\mathbb{N}}$. For all $x\in\mathcal{X}$ it holds
$\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})\right](x)\geq 0$ (B.5)
@@
$J(\mathsf{P}_{t+1})-J(\mathsf{P}_{t})\geq 0$ (B.6)
i.e. the objective function is always increased by a mirror descent iteration. Further, if $\tilde{q}(\mathsf{P}_{t})\in B_{b}(\Omega)$ is such that $\left\|{q(\mathsf{P}_{t})-\tilde{q}(\mathsf{P}_{t})}\right\|_{\infty}\leq\varepsilon_{t}$, then B.5 holds inexactly on $\tilde{q}(\mathsf{P}_{t})$ as
$\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})\tilde{q}(\mathsf{P}_{t})\right](x)\geq-2\varepsilon_{t}.$ (B.7)
@@
$\eta\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})\right](x)\geq D(\pi_{t}(x);\pi_{t+1}(x))\geq 0,$
giving B.5. Integrating B.5 over $\left(\mathsf{Id}-\gamma\mathsf{P}_{t+1}\mathsf{T}\right)^{-*}\nu$ and using the Performance Difference Lemma A.3 one gets B.6. Finally, we get B.7 from
$\begin{split}\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})\tilde{q}(\mathsf{P}_{t})\right](x)&=\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})\right](x)+\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})(\tilde{q}(\mathsf{P}_{t})-q(\mathsf{P}_{t}))\right](x)\\ &\geq\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})(\tilde{q}(\mathsf{P}_{t})-q(\mathsf{P}_{t}))\right](x)\\ &\geq-\bigl{\|}{(\mathsf{P}_{t+1}-\mathsf{P}_{t})(\tilde{q}(\mathsf{P}_{t})-q(\mathsf{P}_{t}))}\bigr{\|}_{\infty}\\ &\geq-\bigl{\|}{\mathsf{P}_{t+1}-\mathsf{P}_{t}}\bigr{\|}\bigl{\|}{\tilde{q}(\mathsf{P}_{t})-q(\mathsf{P}_{t})}\bigr{\|}_{\infty}\\ &\geq-2\varepsilon_{t}.\end{split}$ (B.8)
Where the first inequality follows from B.5, and the latter from the fact that policy operators are Markov operators and have norm 1, and $\bigl{\|}{\mathsf{P}_{t+1}-\mathsf{P}_{t}}\bigr{\|}\leq\bigl{\|}{\mathsf{P}_{t+1}}\bigr{\|}+\bigl{\|}{\mathsf{P}_{t}}\bigr{\|}=2$. ∎